# 111$\quad$1

### v1

@@ 4.3 Digital Human
Based on the current progress of 3D generation, preliminary exploration has also been performed on 4D generation with 3DGS representation. AYG [178] endows 3DGS with dynamics with a deformation network for the text-to-4D generation. It is divided into two stages, static 3DGS generation with the SDS losses based on Stable Diffusion [93] and MVDream [162], and then dynamic generation with a video SDS loss based on a text-to-video diffusion model [189]. In the dynamic generation stage, only the deformation field network is optimized, and some frames are randomly selected to add image-based SDS to ensure generation quality. The generation results are shown in Fig. 13. DreamGaussian4D [25] achieves 4D generation given a reference image. A static 3DGS is first generated using an improved version of DreamGaussian [18]. The off-the-shelf Stable Diffusion Video is utilized to generate a video from the given image. Then the dynamic generation is also realized by the optimization of a deformation network added to the static 3DGS, and the generated video is used as supervision, along with a 3D SDS loss based on Zero-1-to-3 XL [182] from sampled views. Finally, this method also extracts a mesh sequence and optimizes the texture with an image-to-video diffusion model. Last, for video-to-4D generation, 4DGen [190] and Efficient4D [191] both propose utilizing SyncDreamer [192] to generate multi-view images from the input frames as pseudo ground truth to train a dynamic 3DGS. The former introduces HexPlane [50] as the dynamic representation and constructs point clouds using generated multi-view images as 3D deformation pseudo ground truth. The latter directly converts 3D Gaussian into 4D Gaussian and enhances the temporal continuity of SyncDreamer [192] by fusing spatial volumes at adjacent timestamps, achieving time synchronization to generate better cross-time multi-view images for supervision.
@@ 5 Conclusions and Discussions
This survey presents an overview of the recent 3D Gaussian Splatting (3DGS) technique, which not only illustrates how it originates from traditional point-based rendering methods but also how its fast rendering and explicit geometry of the 3DGS representation facilitate a series of works targeting applications like 3D reconstruction and 3D editing. Although 3DGS has greatly improved the efficiency and results’ quality on different tasks, there remain challenges for the 3D Gaussian Splatting, which can be improved by future research.
Robust novel view synthesis
@@
Although 3D Gaussian Splatting has achieved realistic novel view synthesis results, its reconstruction quality degrades as indicated by [35] when dealing with challenging inputs like sparse-view inputs, complex shading effects, and large-scale scenes. Though attempts [68, 22, 79] have been made to get better results, there are still improvement spaces. How to improve its reconstruction robustness on different inputs is an important problem.
Geometry reconstruction.
@@
Despite the efforts on the rendering quality side, few methods [114, 20] work on geometry/surface reconstruction with the 3DGS representation. Compared to the continuous implicit representations like NeRF and SDF, 3DGS’s geometry quality still suffers from its discrete geometry representation. Building up high-quality geometry/surface with the 3D Gaussian Splatting representation would facilitate downstream applications like automated vehicles and animations.
Independent and efficient 3D editing
@@
A few methods have dived into the field of editing 3D Gaussian Splatting’s geometry [29, 26, 28, 20], texture [89, 92], and lighting [22, 95, 21, 96]. However, they cannot decompose geometry, texture, and lighting accurately or need re-optimization of Gaussians’ attributes. As a result, these methods still miss independent editing capabilities or lack efficiency in the editing process. It is promising to extract geometry, texture, and lighting with more advanced rendering techniques to facilitate independent editing and build the connection between 3DGS and mesh-based representation to enable efficient editing.
Realistic 4D generation
@@
With the help of SDS loss based on SD [93], generative models [24, 188, 18] with the 3DGS representation have produced faithful results. However, 4D generation results by current methods [189, 25, 190] still miss realistic geometry, appearance, and physics-aware motion. Integrating data prior like results produced by video generative models and physical laws might boost the quality of generated 4D content.
Platform.
@@
Most implementation of methods based on the 3D Gaussian Splatting representation is written in Python with the cuda-supported PyTorch [193] framework, which may limit its future applicability on wider platforms. Reproducing it with deep learning frameworks like Tensorflow [194] and Jittor [195] can facilitate its usage on other hardware.

### v2

@@ 4.3 Digital Human
Based on the current progress of 3D generation, preliminary exploration has also been performed on 4D generation with 3DGS representation. AYG [264] endows 3DGS with dynamics with a deformation network for the text-to-4D generation. It is divided into two stages, static 3DGS generation with the SDS losses based on Stable Diffusion [140] and MVDream [246], and then dynamic generation with a video SDS loss based on a text-to-video diffusion model [289]. In the dynamic generation stage, only the deformation field network is optimized, and some frames are randomly selected to add image-based SDS to ensure generation quality. The generation results are shown in Fig. 13. DreamGaussian4D [27] achieves 4D generation given a reference image. A static 3DGS is first generated using an improved version of DreamGaussian [20]. The off-the-shelf Stable Diffusion Video is utilized to generate a video from the given image. Then the dynamic generation is also realized by the optimization of a deformation network added to the static 3DGS, and the generated video is used as supervision, along with a 3D SDS loss based on Zero-1-to-3 XL [272] from sampled views. Finally, this method also extracts a mesh sequence and optimizes the texture with an image-to-video diffusion model. Last, for video-to-4D generation, 4DGen [290] and Efficient4D [291] both propose utilizing SyncDreamer [292] to generate multi-view images from the input frames as pseudo ground truth to train a dynamic 3DGS. The former introduces HexPlane [68] as the dynamic representation and constructs point clouds using generated multi-view images as 3D deformation pseudo ground truth. The latter directly converts 3D Gaussian into 4D Gaussian and enhances the temporal continuity of SyncDreamer [292] by fusing spatial volumes at adjacent timestamps, achieving time synchronization to generate better cross-time multi-view images for supervision. SC4D [293] migrates the idea of sparse control point in SC-GS [28] to model deformation and appearance more efficiently. STAG4D [294] proposes a temporally consistent multi-view diffusion model to generate multi-view videos for 4D reconstruction from monocular video or 4D generation. To overcome previous 4D generation models’ object-centric limitation, Comp4D [295] first generates individual 4D objects and later composes them under trajectory constraints. To enable more realistic 3D/4D generation, most methods utilize the priors from diffusion models, which operate on 2D images and require a rendered viewpoint of the generated object. The fast rasterization-based rendering in 3DGS makes the priors more efficiently applied compared to NeRF-based methods with slow rendering speed.
@@ 5 Conclusions and Discussions
This survey presents an overview of the recent 3D Gaussian Splatting (3DGS) technique, which not only illustrates how it originates from traditional point-based rendering methods but also how its fast rendering and explicit geometry facilitate a series of representative works targeting different tasks like 3D reconstruction and 3D editing as shown in the tree diagram in Fig. 14.
Figure 14: Representative works based on the 3DGS representation across different tasks.
Although 3DGS has greatly improved the efficiency and results’ quality on a few tasks, it is not a perfect 3D representation that can satisfy the need of all tasks. Here we first discuss the advantages and disadvantages of commonly used 3D representations including mesh, SDF, NeRF, and 3DGS.
Mesh. Mesh is made of a set of vertices, edges, and faces, which can express detailed geometry at relatively lower storage cost. As the most widely used 3D representation in the industrial, it can create high-quality visual effects with the help of physically-based materials. in real-time However, most meshes are made by artists or creators, which takes a lot of time. Even if some works [296, 297, 298, 299] attempt to generate meshes automatically with neural networks, their generation capability is still constrained by existing datasets with limited quantity.
SDF and NeRF. Both SDF and NeRF are based on the implicit neural fields, which can be learned from a set of multi-view images automatically. The explicit geometry or mesh can also be extracted from SDF or NeRF with the marching cube algorithm. So SDF or NeRF has an advantage in tasks like inverse rendering that rely on a good surface representation. However, due to the dense sampling in 3D space, its rendering is less efficient and limits its applicability on consumer-level devices. Moreover, SDF and NeRF are less successful in dynamic scene reconstruction due to their implicit representation.
3DGS. 3DGS also has explicit geometry but different from mesh it has no edges or faces that connect different Gaussians. To make up for the missing connection information, each Gaussian has anisotropic scales to fill the gap between neighboring Gaussians and synthesize realistic novel views. With the rasterization-based renderer, 3DGS allows real-time visualization of 3D scenes on consumer-level devices, which makes applications like large-scene reconstruction, SLAM, and generation that have efficiency requirements more applicable. Apart from the efficient rendering, its explicit geometry representation enables flexible point reprojection from one viewpoint to another viewpoint, making simultaneous geometry reconstruction and camera pose optimization and dynamic reconstrution easier and tasks like SLAM and large scene reconstruction more efficient. However, due to the discrete geometry representation, the geometry reconstruction quality of current 3DGS-based methods is only comparable to previous SDF-based methods like NeuS [170]. It would be promising to combine other 3D representations with the 3DGS similar to [31, 174] to build up high-quality geometry/surface, facilitating downstream applications like automated vehicles and animations.
After discussing the advantages and disadvantages of different 3D representations, we summarize the challenges that remain for for the 3D Gaussian Splatting and how these challenges might be resolved in the future.
@@
Robust and generalizable novel view synthesis. Although 3D Gaussian Splatting has achieved realistic novel view synthesis results, its reconstruction quality degrades as indicated by [40] when dealing with challenging inputs like sparse-view inputs, complex shading effects, and large-scale scenes. Though attempts [94, 24, 108] have been made to get better results, there are still improvement spaces. How to improve its reconstruction robustness on different inputs is an important problem. In addition, developing a generalizable reconstruction pipeline with or without data prior like [100, 103, 278] would significantly reduce the training cost.
@@
Geometry reconstruction. Despite the efforts on the rendering quality side, few methods [174, 22] work on geometry/surface reconstruction with the 3DGS representation. Compared to the continuous implicit representations like NeRF and SDF, 3DGS’s geometry quality still suffers from its discrete geometry representation.
@@
Independent and efficient 3D editing. A few methods have dived into the field of editing 3D Gaussian Splatting’s geometry [31, 28, 30, 22], texture [134, 139], and lighting [24, 151, 23, 152]. However, they cannot decompose geometry, texture, and lighting accurately or need re-optimization of Gaussians’ attributes. As a result, these methods still miss independent editing capabilities or lack efficiency in the editing process. It is promising to extract geometry, texture, and lighting with more advanced rendering techniques to facilitate independent editing and build the connection between 3DGS and mesh-based representation to enable efficient editing.
@@
Realistic 4D generation. With the help of SDS loss based on SD [140], generative models [26, 277, 20] with the 3DGS representation have produced faithful results. However, 4D generation results by current methods [289, 27, 290] still miss realistic geometry, appearance, and physics-aware motion. Integrating data prior like results produced by video generative models and physical laws might boost the quality of generated 4D content.
@@
Platform. Most implementations of methods and frameworks like GauStudio [300] for the 3D Gaussian Splatting representation are written in Python with the cuda-supported PyTorch [301] framework, which may limit its future applicability on wider platforms. Reproducing it with deep learning frameworks like Tensorflow [302] and Jittor [303] can facilitate its usage on other platforms.